PHARMACEUTICS CLASS

Question:

The activity of sodium ions in a 0.10 M solution is 0.079M. What percentage of ions in this solution is free? Write numerical answer only; round to nearest percentage.
Answer #1

In concentrated ionic solutions, ionic activity is a measure of effective concentration of an ion. Here effective concentration means, the concentration of ions that are actually free and not surrounded by water and opposite ions.

Ionic strength of Na+(aq), I= 1/2 * C*Z2

=> I = 1/2 * 0.10 M * (+1)2 = 0.05 M

At 0.05 M ionic strength, activity coefficient() of Na+(aq) = 0.81 [ From Debye Huckel equation]

Also Activity(a) = activity coefficient() * Free ion concentration(CF)

=> 0.079 M = 0.81 * Free ion concentration(CF)

=> Free ion concentration(CF) = 0.079 / 0.81 = 0.09753

=> Percentage of free ions = (0.09753 / 0.1)*100 = 97.53 %